OpenStudy (anonymous):

2x+3/4=x/4 solve for x

OpenStudy (anonymous):

2x+3/4=x/4 8x+3=x 8x-x=-3 7x=-3 x=-3/7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

here is another one